# MAPS: A Multilingual Benchmark for Global Agent Performance and Security

### v1

@@ 3.1 Translation Pipeline
To ensure semantic integrity, we apply a second binary check: $I(s,\Phi_{\text{enhancement}})\rightarrow{0,1}$. This integrity check detects common LLM failure modes, such as hallucinations, omissions, misinterpretations (e.g., answering instead of translating), and semantic drift. If this verification fails, we revert to the original machine translation (which passed the initial verification test).
The added conditional steps form a robust decision framework: If machine translation is rejected, we use a direct LLM translation; if accepted but enhancement fails integrity verification, we fall back to the machine translation; Otherwise, we use the enhanced translation. This structure ensures graceful degradation, favoring conservative outputs when refinement is unreliable. Formally, the final translation is given by:
$T(s,L_{t})=\begin{cases}\Phi_{\text{direct}}(s,L_{t}),&\text{if }A(s,M(s,L_{t}),L_{t})=0\\ M(s,L_{t}),&\text{if }A(s,M(s,L_{t}),L_{t})=1\text{ and }I(s,\Phi_{\text{enhancement}}(s,M(s,L_{t}),L_{t}))=0\\ \Phi_{\text{enhancement}}(s,M(s,L_{t}),L_{t}),&\text{otherwise}\end{cases}.$ (1)
To ensure the reliability of this pipeline across languages and task types, we conducted human verification on a representative subset of translations. Evaluation design and results are detailed in Subsection 3.3, with additional implementation details in the Supplementary Material.
3.2 Dataset Selection and Composition
Dataset Selection. To support robust multilingual evaluation across agentic capabilities, we construct MAPS benchmark suite based on established agentic AI benchmarks. Those were selected based on four criteria: (1) strong adoption and recognition within the research community, including prior use in agentic evaluation; (2) clearly defined, closed-form answers to enable controlled evaluation; (3) sufficient difficulty to challenge current open-source agents without saturating performance; and (4) practical solvability, ensuring that multilingual degradation can be meaningfully measured. Based on these criteria, we selected four datasets spanning real-world reasoning, software engineering, mathematical problem solving, and security assessment.
SWE-bench. SWE-bench (Jimenez et al., 2023) is a software engineering benchmark constructed from real GitHub issues and associated pull requests across popular Python repositories. Each task presents a bug report and a codebase snapshot, and requires the agent to evaluate whether a proposed patch correctly resolves the issue. We adopt the verified subset [^3] https://openai.com/index/introducing-swe-bench-verified/, in which agents are tasked with validating a patch rather than generating one.
Translated Languages. We selected the following ten typologically and geographically diverse languages: German, Spanish, Portuguese (Brazil), Japanese, Russian, Italian, Arabic, Hebrew, Korean, and Hindi. This selection enables the evaluation of agent performance across a wide range of scripts, linguistic structures, and regional user populations.
Dataset Handling. To preserve the integrity and utility of the original datasets, we applied only minimal and targeted modifications. Across all datasets, we appended translations without modifying or removing any original metadata (such as task type, difficulty level, tools available, etc). Domain-specific syntax—such as equations in MATH, code snippets in SWE-bench, and adversarial prompts in ASB—was preserved exactly, maintaining the original task structure and technical fidelity. For MATH and SWE-bench, which were not originally designed for agentic evaluation, we further applied selective filtering to retain only the most challenging tasks based on the task difficulty field. This follows common practice in prior work to align non-agentic datasets with agentic evaluation settings (Wu et al., 2023), by ensuring meaningful evaluation of agent behavior while avoiding trivial cases.
Data Volume. To balance performance and security evaluation, our benchmark comprises $805$ tasks: $405$ from performance-oriented datasets (GAIA, SWE-bench, MATH) and $400$ from the Agent Security Benchmark. We selected $165$ tasks from GAIA (full validation set), $140$ high-difficulty tasks from MATH ($20$ per topic across $7$ topics), and $100$ hard and medium tasks from SWE-bench. The remaining $400$ tasks include all security-relevant prompts from ASB. Each task was translated into $10$ target languages, and combined with the original English version, this results in a total of $8,855$ multilingual tasks across $11$ languages.
3.3 Translation Implementation and Verification
Translation Implementation Details. We implemented the hybrid translation pipeline described in Section 3.1 using a combination of commercial and open-source tools. For machine translation, we used the Google Translate NMT API  [^4] https://cloud.google.com/translate/docs/advanced/translating-text-v3#translating_input_strings, selected for its support across all ten target languages. To preserve task fidelity, structural elements (e.g., equations, variables, code) that MT systems often mistranslate were temporarily masked and restored after translation. For LLM-based refinement and quality control, we used Cohere Command A and GPT-4o, both multilingual models executed with deterministic decoding (temperature set to zero) to ensure output consistency. System prompts were crafted individually for each task to accommodate domain-specific structures (e.g., code snippets, equations, web URLs), ensuring that the models preserved both intent and format. The code is available  [^5] https://github.com/omerhof-fujitsu/hybrid_translation_demo and representative examples of these prompts are provided in the Supplementary Material.
[…]
Evaluation results confirm high translation quality across the benchmark, with an answerability rate of $94.4$%, corresponding to a total error rate of $5.6\%$. Translations also received average scores of $4.47$ for adequacy, $4.60$ for fluency, and $4.75$ for formatting accuracy (on a 1–5 Likert scale), supporting the benchmark’s preservation of semantic fidelity, linguistic naturalness, and structural integrity. Full per-language results and analysis are included in the Supplementary Material. To support high-precision use cases, we also release a “verified” [^6] https://huggingface.co/datasets/Fujitsu-FRE/MAPS_Verified subset of the benchmark, consisting of 190190~{}190190 translations per language that passed human review across all four datasets.
4 Experiments
We now demonstrate MAPS utility through multilingual evaluation of leading open-source agents.
4.1 Experimental Settings
Agent Assignment per Dataset. To demonstrate the utility of our benchmark, we evaluate open-source agents on each dataset and assess their performance and robustness under multilingual settings. While a unified agent would offer a more broad coverage and controlled evaluation, current systems lack the generalization needed to perform well across diverse tasks (Gioacchini et al., 2024; Chang et al., 2024). To isolate the effect of language variation, we retain each agent’s original configuration, including tools, prompts, and system settings, without any modification.
For GAIA, we used the OpenDeepResearch agent (von Platen et al., 2024), which integrates retrieval, web browsing, and tool use to support real-world reasoning. For MATH, we adopted MathChat (Wu et al., 2024), a zero-shot agent combining multi-turn reasoning with Python execution and the Wolfram Alpha tool. For SWE-bench, we applied SWE-agent (Yang et al., 2024b), which enables autonomous software reasoning through repository navigation, file editing, and test execution. For ASB, we built on the authors’ existing infrastructure and evaluated the original ten-agent system against both direct and indirect prompt injection attacks across a variety of tasks and languages. Each agent was executed using the LLM backbone reported in its original implementation, all of which are considered multilingual models. Specifically, OpenDeepResearch used GPT o1, MATHChat used GPT-4, SWE-bench used GPT-4.1, and ASB used Qwen2. Full configuration details, including model versions and API providers, are provided in the Supplementary Materials.
Experiment Protocol. For each benchmark, the agent was evaluated three times in each of the $11$ target languages, yielding a total of $33$ runs per dataset. We report the mean and standard deviation over these runs. We used the original English task definitions and their translations, without modifying or translating internal agent logic and processing flows like system prompts or tools.
Metrics. We adopt the original evaluation metrics from each benchmark to ensure consistency with prior agent evaluations. For MathChat (Math), we report answer accuracy. For OpenDeepResearch (GAIA), we report the percentage of answers matching either the English or translated reference. For SWE-agent (SWE-bench), we report the percentage of resolved instances, defined as the percentage of submitted patches that successfully resolve the coding issue. For the ASB agent, we report the attack success rate (ASR), a standard metric in the security domain that represents the percentage of attacks that bypass the safety mechanisms. Additionally, we introduce a new metric: Multilingual Effect, which quantifies the performance or security gap between English and the average of all other languages. Given an evaluation metric $M$, the Multilingual effect is defined as follows:
Multilingual Effect =$\text{Multilingual Effect }=\frac{1}{n}\sum_{i=1}^{n}M_{\text{lang}_{i}}-M_{\text{en}}$ (2)
Where $M_{en}$ denotes the performance in English, n is the number of non-English languages in the dataset (in our case $n=10$), and $M_{\text{lang}_{i}}$ represents the performance in the i-th non-English language.
4.2 Results
@@
Figure 4: a) Multilingual effect as a function of the proportion of translated language tokens in input prompts. Each point represents a benchmark-agent pair, with the multilingual effect computed as the average relative degradation in performance or security across non-English languages. The trend suggests a correlation between input translation extent and multilingual vulnerability. b) Relative performance differences from English for each language, broken down by dataset. Negative values indicate a drop in performance compared to English, while positive values (notably in ASB) represent increased vulnerability. The trend highlights how multilingual effects vary by language and task type.
Figure 3 presents the performance of open-source agents across all four benchmarks in English and the ten target languages. In GAIA and ASB, we observe clear performance and security drops: non-English languages consistently underperformed compared to English, with reductions of up to 16% (GAIA) and a rise in vulnerability by up to 17% in ASB. Notably, SWE-bench and MATH exhibit only minor variation across languages, with most scores clustering around the English baseline.
These results reveal important differences in how multilingual degradation manifests across task types. Although all tasks require complex reasoning, some are more constrained than others. For instance, SWE-bench is limited to well-structured Python patches designed to fix specific test cases. As a result, the reliance on natural language explanations is reduced, with greater emphasis placed on strict Pythonic syntax and code correctness. On the other hand, GAIA focuses on solving real-world tasks with much more flexibility. Thus, the importance of the natural language problem statement is significantly higher. Additionally, in benchmarks like MATH and SWE-bench, the opportunity for translation is inherently limited, as a large portion of the input consists of mathematical notation or source code, thus naturally reducing the multilinguality effect. To understand this variation, we examine a potential driver: the proportion of localized, target-language-oriented input in each benchmark in each benchmark’s input.
Interestingly, we observe that Japanese yields the lowest ASR (attack success rate) in the ASB benchmark, indicating the highest robustness to adversarial inputs. This result can be partially attributed to the fact that the ASB agent was implemented using the Qwen2 model  (Yang et al., 2024a), which is known for its strong alignment for Japanese language tasks. Qwen2 has consistently demonstrated strong performance in Japanese-specific LLM benchmarks and leaderboards [^7] https://rinnakk.github.io/research/benchmarks/lm/index.html, suggesting that alignment and fine-tuning in a particular language can significantly enhance resilience against multilingual adversarial prompts. This reinforces the importance of language-specific alignment training in the development of robust and secure agentic systems.
Figure 4 examines the relationship between prompt composition and multilingual performance. Part (a) shows a correlation between the percentage of non-English tokens in the input and the average performance gap (relative to English) across all four datasets. Benchmarks with higher proportions of localized, target-language-oriented input, such as GAIA and ASB, exhibit greater degradation, whereas SWE-bench, with predominantly English input (e.g., code), shows higher preservation.

### v2

@@ A.3 Hybrid Translation Pipeline Implementation Additional Information
To ensure semantic integrity, we apply a second binary check: $I(s,\Phi_{\text{enhancement}})\rightarrow{0,1}$, targeting common LLM failure modes, such as hallucinations, omissions, misinterpretations (e.g., answering instead of translating), and semantic drift. If this verification fails, we revert to the original MT output (which passed the initial verification test).
These conditional steps form a robust decision framework: If MT is rejected, we use a direct LLM translation; if accepted but enhancement fails integrity verification, we fall back to the MT; Otherwise, we use the enhanced translation. Importantly, since both MT and LLM-based translation are robust technologies, we did not encounter any scenario in which both systems failed to produce a valid translation. A formal expression of this pipeline is provided in the Appendix.
For the translation pipeline, we used Python 3.11 alongside a hybrid setup comprising Google’s Neural Machine Translation (NMT) API, Cohere Command-A, and GPT-4o. To support high-quality multilingual alignment, we crafted dedicated system prompts for each stage of the pipeline—translation verification, semantic enhancement, and structural integrity checks—tailored to the specific requirements of each domain. For example, GAIA tasks involve real-world decision-making, where preserving subtle intent and contextual detail across languages is critical. In contrast, MATH tasks require precise handling of symbolic notation and mathematical terminology, necessitating strict preservation of LaTeX expressions and domain-appropriate language. An example of a task-specific system prompt is provided in Fig. 6. The full translation pipeline implementation, including prompt templates and evaluation logic, is available  [^4] https://tinyurl.com/hybrid-translation-pipeline.
To evaluate translation quality across different configurations, we employed LLM-based evaluation using a structured prompt. This approach follows the framework introduced by Jiang et al.(Kocmi and Federmann, 2023), who show that LLMs can reliably assess translation quality when guided by clear criteria. We adapted and extended their methodology to include four dimensions tailored to our benchmark context (those metrics are similar to the ones we define in the manual verification experiment): (1) adequacy—semantic fidelity to the original task, (2) fluency—naturalness and grammaticality in the target language, (3) formatting accuracy—preservation of structural elements such as LaTeX, code, or tool references, and (4) answerability—whether the translation remains usable and complete. The evaluation prompt used is provided in Fig.7. Unlike traditional evaluation metrics such as COMET (Rei et al., 2020), which rely on the same-language reference translations and embedding-based comparisons, our agentic tasks often lack ground truth references and span cross-language instruction. In such settings, reference-based metrics become inapplicable, whereas LLM-based zero-shot evaluation enables context-sensitive, reference-free scoring.
Using the LLM as a judge approach, we evaluated four translation strategies (Table 4) by translating the GAIA validation set, to empirically identify the most effective pipeline design. This aligns with the analysis in Section 3 of the main manuscript, which discusses the trade-offs between machine translation (MT) and LLM-based generation. As expected, Google NMT showed strong performance on formatting but slightly lagged in semantic adequacy and fluency. Command-A improved fluency and meaning preservation but occasionally introduced structural drift. Hybrid approaches—where NMT outputs were refined by either GPT-4o or Command-A—consistently outperformed single-method setups. Notably, the Google+Command-A variant achieved the highest scores across all metrics, including adequacy (4.76), fluency (4.90), formatting accuracy (4.96), and answerability (95.9%). These results validate our decision to adopt a hybrid approach and extend prior work(Ki and Carpuat, 2024) by integrating domain-specific prompting, automated checks, and structure-aware refinement for multilingual translation in agentic AI benchmarks.
Translation Method Adequacy (1-5) Fluency (1-5) Formatting Accuracy (1-5) Answerable (%)
Google Translate NMT 4.667 4.870 4.946 92.9%
Command A (LLM only) 4.704 4.872 4.934 93.9%
Hybrid (Google Translate NMT + GPT-4o) 4.709 4.832 4.923 94.3%
Hybrid (Google Translate NMT + Command A) 4.760 4.899 4.959 95.9%
Table 4: Translation quality comparison across methods based on human evaluation. Bolded values represent the best-performing configuration for each metric.
Figure 7: System prompt used to guide the LLM-based translation evaluation. The prompt instructs the model to assess translations across four dimensions—adequacy, fluency, formatting accuracy, and answerability—and return results in a strict JSON format.
A.4 Hybrid Translation Manual Evaluation Additional Information
To ensure high-fidelity multilingual translation in MAPS, we conducted a human evaluation of translated samples using a standardized annotation protocol. This procedure mirrors the evaluation framework used for our LLM-based assessment, following the methodology proposed by Jiang et al. (Kocmi and Federmann, 2023). Below, we reproduce the instructions provided to annotators, which guided their evaluation of semantic adequacy, linguistic fluency, formatting consistency, and answerability across all benchmark domains.
A.4.1 Translation Verification Guidelines for Annotators
Purpose. You are asked to verify the quality of translations from English into your target language for a multilingual benchmark dataset. Your careful evaluation ensures that the translations are accurate, fluent, and correctly formatted.
Dataset Overview and Translation Challenges. You will review 200 translated samples drawn from four benchmark datasets:
  • GAIA (50 samples): natural language questions requiring web search and file analysis.
  • Agent Security Benchmark (60 samples): tasks probing unsafe or manipulative agent instructions.
  • MATH (70 samples): algebra, calculus, probability, and geometry problems, often containing LaTeX and variables.
  • SWE-Bench Verified (20 samples): GitHub issues involving bug descriptions and code snippets.
Evaluation Procedure. For each sample:
  1. 1.
    Read the English source text.
  2. 2.
    Read the translated text in the target language.
  3. 3.
    Assign the following ratings:
    • Adequacy (1–5): how well the meaning is preserved.
    • Fluency (1–5): how natural and grammatically correct the translation is.
    • Formatting Accuracy (1–5): whether LaTeX, variable names, units, and structure are preserved.
    • Answerable? (Yes/No): Answer No if you feel you cannot answer the original question because of a major translation mistake (e.g., reversed meaning, missing information). Otherwise, mark Yes.
    • Notes (optional): any remarks you wish to add.
Rating Scales.
  • Adequacy: 5 = perfect meaning preservation, 1 = major meaning loss.
  • Fluency: 5 = native-like fluency, 1 = broken or ungrammatical.
  • Formatting Accuracy: 5 = all formatting correct, 1 = major formatting errors.
Important Notes.
  • Do not correct or rewrite the translations.
  • Pay special attention to LaTeX symbols, variable names, and numerical units.
Sheet Columns Example (one row per sample):
  • Item ID – Sample identifier.
  • English Source Text – Original English input.
  • Translated Text – Translated version.
  • Adequacy (1–5) – Meaning preservation.
  • Fluency (1–5) – Naturalness and grammar.
  • Formatting Accuracy (1–5) – Structural fidelity.
  • Answerable? – Unanswerable due to error?
  • Notes (Optional) – Free-text comments.
Figure 8: Translation verification examples submitted by annotators during manual evaluation. Left: A successful Japanese translation from the GAIA dataset, rated highly across all dimensions with only a minor terminology note. Right: A Korean MATH example flagged as critically erroneous due to semantic and grammatical mismatches in imperative phrasing. These examples illustrate the evaluation schema’s ability to surface both subtle formatting issues and major adequacy failures.
A.4.2 Translation Verification Extended Results
To ensure the validity of our multilingual benchmark, human verifiers manually reviewed approximately 25% of the translated data—corresponding to over 2,000 samples—across all tasks and languages. In the main manuscript, we reported the overall scores for adequacy, fluency, formatting accuracy, and the error rate derived from answerability. In Tables 6, 7, 8 and 9 we present the full breakdown of results for each individual dataset and in Table 10 we conclude with the manual verification average results across all datasets. Across all four benchmark datasets, native verifiers’ ratings reveal high translation quality, with average adequacy (4.47), fluency (4.60), and formatting accuracy (4.76) on a 1–5 scale. Notably, answerability remained consistently high (mean: 0.94), suggesting that most translations retained functional fidelity. Formatting consistency was strongest for ASB and GAIA, while fluency slightly declined for Arabic and Japanese, correlating with minor adequacy drops. Additionally, Fig. 8 presents one successful and one failed translation example, as annotated by Japanese and Korean evaluators, respectively. These results confirm that closed-form multilingual alignment is achievable across diverse tasks without compromising validity.
A.4.3 Statistical sufficiency of the 25% manual audit
Let n=2,000n=2{,}000 be the size of our manually verified subset ($\approx 25\%$ of the full benchmark). For any binary outcome—here, Answerable vs. Unanswerable—the sample proportion $\hat{p}$ is an unbiased estimator of the true corpus proportion $p$. Under the binomial model, its standard error is
$\mathrm{SE}(\hat{p})=\sqrt{\frac{\hat{p}(1-\hat{p})}{n}}.$
The variance is maximised when $\hat{p}=0.5$, giving a worst–case standard error $\mathrm{SE}_{\max}=0.5/\sqrt{n}$. Using the normal approximation with critical value $z_{0.025}=1.96$, the 95 % two–sided margin of error is therefore
$E=1.96\times\frac{0.5}{\sqrt{n}}=\frac{0.98}{\sqrt{n}}=\frac{0.98}{\sqrt{2{,}000}}\approx 0.022.$
Hence any proportion reported from the manual audit has a 95 % confidence interval of at most $\pm 2.2$ percentage points. Put differently, any systematic translation error affecting more than $\approx 2\%$ of the corpus would almost certainly have been detected in our audit. Because the sample was stratified by language and dataset, the same formula applies within each stratum with $n_{\ell}\!=\!0.25\,N_{\ell}$ items audited per language $\ell$. Putting the numbers together, the overall Answerable estimate $\hat{p}=0.944$ therefore has a 95 % confidence interval
$0.944\pm 0.022\;=\;[0.922,\;0.966].$
Thus, even in the worst case, the corpus-level answerability remains above 92 %.
Language Variant
Chinese Simplified (Mainland)
Russian Standard (Russia)
Arabic Modern Standard Arabic
Hebrew Modern Israeli
German Standard German
Spanish Latin American
Hindi Standard Hindi
Japanese Standard Japanese
Korean Standard Korean
Italian Standard Italian
Portuguese Brazilian Portuguese
Table 5: Dialect used per language.
GAIA
Lang Adequacy (1–5) Fluency (1–5) Formatting Accuracy (1–5) Answerable
de 4.14 4.04 4.74 0.91
ja 4.30 4.70 4.91 0.88
he 4.20 4.62 4.86 0.88
ko 4.49 4.84 4.40 0.94
pt-br 3.93 4.16 4.53 0.96
it 4.26 4.24 4.72 0.91
es 4.18 4.24 4.99 0.94
ar 3.54 3.79 4.41 0.98
hi 4.22 4.52 4.72 0.92
ru 4.84 4.99 4.90 0.96
AVG 4.21 4.41 4.72 0.93
Table 6: Manual translation evaluation scores for the GAIA dataset across 10 languages. Metrics include adequacy, fluency, formatting accuracy (all rated on a 1–5 scale), and answerability (proportion of translations judged as answerable).
MATH
Lang Adequacy (1–5) Fluency (1–5) Formatting Accuracy (1–5) Answerable
de 4.80 4.87 4.97 1.00
ja 4.11 4.41 4.96 0.82
he 4.61 4.46 4.79 0.94
ko 4.62 4.95 4.66 0.93
pt-br 4.95 4.84 4.62 1.00
it 4.65 4.85 4.88 1.00
es 4.74 4.49 4.28 0.87
ar 3.78 3.99 3.81 0.82
hi 4.76 4.79 4.96 0.95
ru 4.75 5.00 4.75 0.99
AVG 4.58 4.67 4.67 0.93
Table 7: Manual translation evaluation scores for the SWE-Bench dataset across 10 languages.
SWE-Bench
Lang Adequacy (1–5) Fluency (1–5) Formatting Accuracy (1–5) Answerable
de 4.30 4.80 4.50 1.00
ja 4.40 4.05 5.00 0.85
he 4.55 4.50 4.50 0.90
ko 4.35 4.25 4.40 0.90
pt-br 4.75 4.80 4.75 1.00
it 4.55 4.65 4.35 0.95
es 4.00 4.50 4.50 0.95
ar 3.60 4.30 4.20 0.90
hi 4.40 4.35 4.80 0.90
ru 4.85 4.95 4.80 0.95
AVG 4.375 4.515 4.580 0.930
Table 8: Manual translation evaluation scores for the SWE-Bench dataset across 10 languages.
ASB
Lang Adequacy (1–5) Fluency (1–5) Formatting Accuracy (1–5) Answerable
de 4.10 4.53 4.97 0.95
ja 4.68 4.76 4.91 0.96
he 4.68 4.67 5.00 0.92
ko 4.95 4.93 4.96 0.98
pt-br 4.72 4.55 4.92 1.00
it 4.25 4.47 4.80 0.98
es 4.57 4.40 4.87 0.97
ar 4.45 4.68 4.90 0.98
hi 4.88 4.85 4.95 0.98
ru 4.82 4.90 4.83 0.98
AVG 4.61 4.67 4.91 0.97
Table 9: Manual translation evaluation scores for the ASB dataset across 10 languages.
All Benchmarks (Average)
Lang Adequacy (1–5) Fluency (1–5) Formatting Accuracy (1–5) Answerable
de 4.375 4.554 4.867 0.964
ja 4.356 4.553 4.936 0.881
he 4.525 4.565 4.840 0.915
ko 4.660 4.847 4.660 0.943
pt-br 4.609 4.580 4.703 0.989
it 4.422 4.565 4.763 0.968
es 4.475 4.403 4.657 0.922
ar 3.859 4.255 4.423 0.936
hi 4.625 4.697 4.881 0.947
ru 4.803 4.961 4.817 0.974
AVG 4.471 4.598 4.755 0.944
Table 10: Average manual translation evaluation scores across all benchmarks (GAIA, MATH, SWE-Bench, and ASB). Metrics include adequacy, fluency, and formatting accuracy (rated on a 1–5 scale), and answerability (proportion of samples deemed answerable).
A.4.4 Language Variants Used
To ensure wide applicability across regions, we translated the dataset into 11 languages using the most widely recognized and formal variant in each case. The table below lists the dialect or regional standard used in the translation process.
Figure 9: Trajectory variability between English and Hindi agents in solving the same question. The English agent follows a computation-driven path—extracting raw values and calculating the result—whereas the Hindi agent relies on surface-level textual matching and assumed equivalence. This leads to reasoning failure due to the absence of explicitly stated percentages in the Hindi query context.
@@ A.4.4 Language Variants Used
Figure 10: Relative performance differences from English for each language, broken down by dataset. Negative values indicate a drop in performance compared to English, while positive values (notably in ASB) represent increased vulnerability. The trend highlights how multilingual effects vary by language and task type.
A.5 Experimental Settings Additional Information
We conducted all experiments on a Linux-based machine (Ubuntu 22.04) equipped with 64 virtual CPUs and 256 GB of RAM. To ensure reproducibility and compatibility, each agent was executed within its own isolated software environment tailored to its specific dependencies. To ensure a fair evaluation of multilingual effects (see Eq. 2 in the main manuscript), we preserved each agent’s original implementation without introducing any code modifications. This isolation strategy allowed us to assess the impact of language variation on agent performance and behavior without confounding factors from architectural or behavioral changes. Datasets were also left unchanged in structure, format, and content. Only the task inputs and expected outputs were translated into the target languages. This process yielded 12 language-specific evaluation suites per dataset (including English), each of which can be executed independently, enabling controlled multilingual evaluation with minimal setup and without requiring code adaptation.
A.5.1 Agents Configuration Additional Information
For the GAIA benchmark, we used the publicly available agent from OpenDeepResearch (von Platen et al., 2024), which features a modular framework supporting document retrieval, browser-based execution, and web search. The agent is designed for general-purpose, real-world reasoning and incorporates memory and planning via tool-augmented prompting. In our evaluation, we disabled the web search functionality due to its reliance on the external Serper API, which imposes rate limits and introduces latency variability. This omission was applied across all languages to ensure consistency, and throughput scalability under multilingual evaluation. For MATH, we adopted MathChat (Wu et al., 2024), a mathematical reasoning agent that combines multi-turn chain-of-thought prompting with symbolic computation using Python and the Wolfram Alpha API tool.
For SWE-bench, we used SWE-agent (Yang et al., 2024b), an autonomous software engineering agent capable of navigating repositories, modifying code, and executing test suites. It employs a plan-execute loop enhanced with file-level attention and dynamic tool invocation. For ASB, we built on the original benchmark infrastructure (Zhang et al., 2024), eliciting synthetic adversarial tool-use in various agentic scenarios (e.g. travel-agent, legal consultant, system admin).
Each of the above was executed using the same LLM backbone reported in its original publication. All selected models are considered to be multilingual models. Specifically, GAIA used GPT-O1 (model version: 2024-12-17), MATHChat used GPT-4 (model version: turbo-2024-04-09), SWE-agent used GPT-4.1 (model version: 2025-04-14), and ASB used Qwen2 7B (model version: 2024-10-01).
A.6 Results Additional Information